OpenStudy (anonymous):

If the discriminant of an equation is zero, which of the following is true of the equation? A. It has one complex solution. B. It has two real solutions. C. It has two complex solutions. D. It has one real solution.

OpenStudy (thomas5267):

Consider the quadratic formula. \[ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}=\frac{-b\pm\sqrt{\Delta}}{2a},\text{ where }\Delta\text{ is the discriminant.} \] What happens when \(\Delta=0\)?